OpenStudy (anonymous):

Solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21

OpenStudy (unklerhaukus):

the difference of the first two equations 5y-4z =-5, y = -1/5 z

OpenStudy (unklerhaukus):

the sum of the last two 2x + 2z = -14, x + z = -7

OpenStudy (anonymous):

x=-2 y=3 z=5

OpenStudy (turingtest):

Or learn linear algebra, it makes this stuff so much easier!